April 1997 Archives

Apr 22 1997

Valery Ryumin
RSC Energiya
4 Lenin Street
Kaliningrad
Moscow Region, 141070
RUSSIA

Greetings, Valeri!

The information you have sent us recently on the Mir systems status, plans for repair, and the consumables status has been very helpful in preparing for the upcoming Flight Readiness Reviews. I would like, However, for us to be able to present a joint picture of the Russian criteria for mission continuation and out rationale for concurrence with that criteria. I am forwarding a sample of some of the areas of concern that we feel should be addressed to answer the questions of senior NASA management. This information is based on the information from you and existing joint documents. Please review this and we will discuss it at our regular telecon on Tuesday. And please understand that this is in addition to current status. We are trying to show that we understand each other's safety criteria and mission management philosophy so that upper management can let us do our jobs without anxiety.

A part of the American approach to managing a spacecraft in flight is to develop a minimum equipment list which identifies the set of hardware that must be operational to initiate or continue the mission. This makes perfect sense when you are preparing to launch a spacecraft such as the Shuttle. Some of our managers expect us to provide such a list indicating what must be operational on the Mir before we would agree to transfer Foale or otherwise continue the mission, We understand that the Russian side does not manage Mir in that fashion, and we are trying to communicate what we believe is your philosophy to our management.

NASA MIR 4 mission, MD 93-96 (Edited)
Date: 14 April 1997 / Mission Day 93
Mir Station System/Subsystem Activities
KOB-2 Loop Leak Repair (entire crew)

Mission Science Payload Activities: QUELD Sample E-13;DOZA - TLD Processing

Stowage/Transfer
The crew reported the failed solid oxygen generator as well as the cut-out segments of the thermal pipes have been packed in ziplock bags in preparation for their return on STS-84.

The crew also mentioned that the 'Raketa' vacuum cleaner they have been using throughout their increment is failing. The crew is unsure if the black old vacuum cleaner they found onboard is functioning. The crew recommended manifesting a replacement ASAP and suggested STS-84. The ground gave no response to this suggestion.

Mir Systems

The Elektron system has been working continuously since Saturday evening. Its O2 production is not quite enough for three crewmembers, so the O2 is being supplemented with Progress O2 as required. This procedure is described in on-board documentation. They hope to maintain this mode of operation until Shuttle arrives with the new Elektron.

The Vosdukh system is working well. It's efficiency (with respect to removing moisture) will be improved by installing a fan to blow some cooling air onto the system.

NASA MIR PROGRAM
Mir 23 / NASA 4 Status
Date: 13 April 1997
Mission Day: Mir 23 / NASA 4 MD 63 / 92

Mir Station System/Subsystem Activities Mir IFM Repair (entire crew) Rest (entire crew)

SMP/Medical: per day according to the first day of the exercise Cyclogram (treadmill and expanders). A radiogram from the Russian medical group has been prepared for uplink today or tomorrow, recommending that the commander and Jerry return to two hours of exercise per day in preparation for the EVA.

Private Family Conference
Successful 11 minute audio PFC between Jerry and Kathryn.

Mir Systems
Friday and Saturday, the Elektron system continued to shut down automatically after runs varying in duration between 3 and 20 minutes. Since last night, the system has run continuously for 22 hours without a problem (as of 1445 DMT). The crew was instructed to stop using air from the Progress, since the Elektron is now producing sufficient oxygen.

The Vosdukh system is removing CO2. The ppCO2 is now about 4.8 mmHg. The system is operating in an `off-nominal' mode. The lack of cooling (due to the VGK shutdown) is causing the systems water vapor removal to be less efficient, resulting in more water than usual being vented overboard with the CO2.

Leak checks continue on the VGK system in Kvant. There is some concern that the leak may be in a hydroconnector at a point with little or no access. More panels will be opened for examination tomorrow.

The Urine Processing System is also operating in an off-nominal mode. A Radiogram has been sent up to the crew with instructions on how to over-ride or work-around a problem with a micro-pump in the system. Urine is being transferred to one of the tanks on the Progress.

NASA MIR PROGRAM
Mir 23 / NASA 4 Status
Date: 12 April 1997
Mission Day: Mir 23 / NASA 4 MD 62 / 91

Mir Station System/Subsystem Activities Elektron maintenance
Condensate Presence Check of Mir (Mir 23) Wet Cleanup (entire crew)

Mission Science Payload Activities
Sleep Study Final Ops and Stowage (NASA 4)BTS Filter Check (NASA 4)

SMP/Medical
The crew requested to have their MK-12, Reflotron blood analysis data voiced up. They also suggested using venous blood instead of capillary blood for the next MK-12 session. The crew feels they would not have any difficulties drawing venous blood. Furthermore, the fingersticks during the last session caused sore finger tips for a couple of days.

During the last comm pass last night, the crew reported exercise was not performed since they were tired.

Mir 23 / NASA 4 Status (edited)
Date: 11 April 1997 / Mission Day: 90
Mir Station System/Subsystem Activities
IFM's (entire crew)

Mission Science Payload Activities
Sleep Study (NASA 4)
CSST (NASA 4)

PAO: Press Conference held today between Jerry and reporters from JSC/KSC. Comm quality was spotty with a late start due to Mir-related discussions and a two-minute comm dropout. It was also terminated two minutes early by the SFD for further systems discussion. The MOST team will discuss the abruptness of the cutoff with Mr. Blagov.

Mir Systems
Per a conversation with V. Blagov, the VGK in Kvant 1 is still leaking, and therefore, will not provide cooling to the Vosdukh. Mr. Blagov said the plans are to operate the Vosdukh without cooling. He also said the Vosdukh system used in the design and ground testing of the BURAN had no thermal interface, so there is no concern with operating the Vosdukh for at least 30 days.

In this mode of operation, the system will dump moisture out of the cabin, so

the 30 day limit is related to the lower limit of ppH2O. At 09:49 DMT the ppH20 was 13.92 mm Hg. The Vosdukh was turned on in this mode before the 11:16 DMT pass, at which the crew reported nominal operations, i.e. no alarms. The nominal temperature limit for the CO2 absorbing beds, without degradation of efficiency, is 23C. If these beds are exposed to 50C air, the system will stop working (the CO2 removal beds will not remove CO2). The plans are to operate the Vosdukh in Kvant 1 in this mode until the leak in VGK can be found and repaired. The crew will be working on this activity until the leak can be found.

The crew is systematically searching for the leak in VGK. This includes isolating segments of lines and monitoring for a pressure drop. The crew was also asked to carefully inspect the hydroconnectors.

The Antares transmitter (which will be ready for shipment to the US on 20th April) will replace one of the two transmitters. One is hard failed, and the other is operational, but because of the lack of cooling (from KOB-1) to it, it will only operate for 10 minutes.

Date: 10 April 1997 / Mission Day 89
Mir Station System/Subsystem Activities: IFM's (entire crew)

Mir Systems

1. Jerry and Vasily (Tsibliev) completed removal of the leaking heat exchanger in Kvant, while Sasha (Lazutkin) worked on the Elektron filter bypass in Kvant II.

2. The pressure check of the Kvant VGK cooling loop showed the pressure decreased from 900 mmHg to 770 mmHg in 30 minutes. By the next pass, the crew had repressurized the loop to 898 mmHg, and by 20:01, had held pressure for over one hour. All heat exchangers are still clean and dry.

3. The Elektron in Kvant II was switched on, but then automatically shut down after 9 minutes. A malfunctioning BVKD pressure gauge was also detected. The crew was told to repeat the IFM procedure. It was suggested a new BVKD pressure gauge be added to the STS-84 manifest. At 20:02 DMT, the crew indicated they had activated the Elektron three additional times per the ground instructions and the unit worked for one minute, two minutes, and twelve minutes, respectively, before going into 'auto shutdown'. The ground said telemetry indicated a bad control unit and instructed the crew to swap the unit with the unit from the Kvant Elektron.

4. The Condensate Removal System (CRS) cooling fan in the Soyuz was showing 'negative function.' Currently, all Mir humidity removal is being performed by the Soyuz CRS. The ground is currently writing an IFM r/g for this.

5. According to the most recent 'Mir Onboard Systems and Consumables Status' received today, there are 14 unused LiOH canisters onboard as of yesterday.

Shuttle Mir Science Program / NASA 4 / Mir 23 Status
Date: April 3, 1997 - April 9, 1997
Mission Days: 82-88 (NASA 4) / 53-59 (Mir 23)

Mir Systems

Telemetry - The OMOTsAO unit failed yesterday; this device stores uplinked executable commands, in particular, the command for the MIPS downlink. That means that there will be no MIPS downlink available until the system is repaired. Repair activities on this unit are scheduled for 14 April.Suggestion will be made to Jerry to use the Packet system. (Jim Medford and Ed Walters (MOD) are trying to understand what the MOTsA is. The SM IPT is going to see if a similar unit is in phase 2. This is the first appearance of this in our documentation.

Progress - Successful Auto-docking of the Progress occurred last night at 20:30 DMT. The crew was directed to open the hatch early last night and to temporarily suspend using candles, utilizing the O2 from the Progress. Not sure if ethylene glycol was carried on this flight or how much in case they need to replace the amount leaked out from the thermal system. They were not sure if another solid oxygen generator unit was flown up. The POSA will get the manifest and send it to me in the next week or so.

I have asked for a contact on understanding how they use the O2 from the Progress to replemish the MIR oxygen. Jim Medford is the contact to get this info. The Russians apparently use this process all the time (ie, not just as a result of the recent failures). I therefor suspect we will do it for Station and suggest it would be helpful to understand this capability. I believe we have maintained a record of the atm content pre and post Progress docking.

Repair plan - Two crew members working on removing the leaking heat exchanger from the VGK loop in Kvant 1. The crew will cut and cap the pipes leading in and out of the leaking heat exchanger in the Kvant 1.After the loop has been leak checked, the Vozdukh system will be reactivated. The goal is to complete this activity and reactivate the loop by the end of the 10th. The ground personnel are confident that the remaining three heat exchangers will be adequate to handle the system load.

The other crew member is working on the Elektron system in Kvant 2. The original goal was to repair and re-start the unit on the 9th. We do not have a status on this yet.

Date: 9 April 1997 / Mission Day: 88

Experiment Operations: Victor Blagov informed us Energia management has set a deadline of the 15th of April to successfully complete the maintenance activities. If these activities are not finished, then the EVA may be in jeopardy. Furthermore, he requested we clear Jerry's calendar of science through Friday, 11 April, to permit him to work on the maintenance activity. As a result, Jerry will perform only the Sleep experiment and CSST on Friday, April 11. Jerry on a later pass said he does not have any problem performing system maintenance activities, but wanted to make sure the cancellation/delay of science is analyzed carefully.

SMP/Medical: The time reserved for a PMC was used for a regular US comm pass on the open air-to-ground. Soloviev spoke with the crew on the open air-to-ground. He made two 'commands'. First priority: the crew obey the work-rest schedule with a full eight hours of sleep each night. Second priority: No science until station repair operations are complete.

Date: 8 April 1997 // Mission Day 87

Mir Systems

1. Progress 234 - At 19:05 DMT, the crew was informed that ground commands to activate the two TORU (manual docking system) transmitters onboard Progress 234 were unsuccessful. The decision was made for one attempt using the Automatic Docking System. By 20:23 DMT the TORU transmitters were reported to be up and functioning.

2. Successful Auto-docking occurred at 20:30 DMT; hatch opening occurred prior to the next comm. pass. Direction for early hatch opening tonight was given to utilize the O2 onboard the Progress. As of 22:30 DMT the crew has not entered the Progress.

3. On the day after docking, two crew members will be scheduled to work on removing the leaking heat exchanger from the VGK loop in Kvant 1. The other will be working on the Elektron system in Kvant 2.

Date: 3 April 1997 / Mission Day 82

Mir Station System/Subsystem Activities
Replace PTAB Unit #2 in Priroda
Clean Mir power connectors x-301 and x-302

Mir Systems

Mr. Blagov reported the Vosdukh (CO2 Removal System) was shutdown due to a lack of thermal control from the VGK system in the core module. Some time ago, the core module's VGK and KOH-1V loops (which were originally independent loops) were re-plumbed to create one loop. This was done to reduce the number of required pump panels (one pump panel can support both loops). Earlier in the week, the pressure of the KOH-1V began dropping. Today, the delta pressure at the pump panel dropped to zero, resulting in a shutdown of the loop. The plans are to separate the loops back to two independent loops and determine if another leak exists and isolate it. Mr. Blagov estimated the on-board supply of LiOH canisters would last 10 days, and stated a 15 day supply is manifested on the Progress, docking next week. Currently, the priority of the loop repairs has not been established.

Note: This VGK loop is a separate loop from the one discussed in earlier reports. At last report, the other VGK loop (in Kvant 2) was being patched with an IFM procedure the ground had sent up, and was nearly complete.

Mir 23 / NASA 4 Daily Status Reports (Edited) // Period of April 2-3, 1997
Date: 2 April 1997 / Mission Day 81

Experiment Hardware:

1. EDLS - Jerry saw a green light flash inside the drive door and on panel every three seconds regardless of activity. We have not pursued this report as a high priority item. At the present time it is believed that all of the MiSDE PCMCIA disks are full. We are waiting to hear Jerry's report on the disk status before determining whether any future EDLS sessions will be scheduled.

2. Orientation - It has been reported by the crew the Orientation session ran into difficulties when Jerry was unable to activate the system. This is still being investigated by the Russian specialists; however, it appears that it is possible, repeat, possible, radiation has hit the PCMCIA cards. They are believed to be static PCMCIA cards of the type originally used by MIPS.

Video System - Video downlink of the CNN Interview was accomplished. The video quality was poor; according to the Russian specialist, the signal was distorted due to problems between Russian ground sites.

Subject: Telecon Notes
Date: Tuesday, April 01, 1997 9:36AM
Regular Tuesday Ops Telecon 4/1/97

1) Shuttle Status - Mentioned RCS seal-saver problem on OV-104, no launch impact expected as of now.

2) Mir Status - Yuri A. reviewed thermal loop status. Due to Kvant II internal loop leak that module is running about 3 deg C higher than normal but still within allowed limits. Also one of the solar array batteries (sets?) is running a little warm due to the reduced cooling. They plan to use the sealant to repair the leaks tomorrw but it will take a total of three days (curing time) before the loop is running. Victor emphasized that this is the only loop they think is leaking into the cabin, the other loop leaks are believed to be internal....

Victor also pointed out that while everybody in Houston is worried about O2 generation, the Russians are feeling far more vulnerable in the thermal loop area.

Repair plans/order after Progress is 1) Elektron filter bypass (3 hrs), 2) thermal loop leak troubleshooting (flourescent dye tracer), and then EVA preps April 15th).

3) They have received the copy of the STS-81 joint incidents report and have several 'updates' TTI/B. Fingerov will provide them with an electronic copy of the TTI translation to facilitate markups.

4) The strategy for the potential addition of the sixth docked day was discussed. Victor proposed that the first activity was to transfer and install the Elektron, on the first docked day, even before seat transfer. That would keep from breaking up the currently planned sequence of docked activities. (I do not see any problems with this, if Phase 1 is OK with it). The existing Elektron unit will already have been removed before docking.

Note: Victor indicated that we were returning the fluid units for BOTH elektrons on Mir now. This is not my understanding. He said the electronic control modules are OK, only the fluid units need to be returned. I thought we have told Vorobiev we did not have room for both. Also, this proposal gives insight into their confidence in getting the onboard unit working....

5) Other topics addressed were STS-86 timelines (sleep cycle) and changes to the comm and interactions document 3245. Victor mention that THEY HAD TALKED TO THE 86 CREW WHILE THEY WERE IN MOSCOW AND TALKED ABOUT ADDING A 6th DOCKED DAY... More work remains on these last items.


Loading

 



Monthly Archives

About this Archive

This page is an archive of entries from April 1997 listed from newest to oldest.

March 1997 is the previous archive.

May 1997 is the next archive.

Find recent content on the main index or look in the archives to find all content.